# On the Spectral Clustering of a Class of Multigrid Preconditioners

### v2

@@ 6 Adaptation of the direct solver to practice
We want to emphasize that, in our opinion, the results of the above sections are important on their own in raising the hypothetical flagpole in the form of Algorithm 2 combined with Corollary 4.5. This gives practitioners a clear message of what are we trying to approximate for the Algorithm 2 setup and what can we expect in return.
@@ 6.1 Conditioning of the preconditioned system
We start by numerically testing the direct solver version of Algorithm 2 with Corollary 4.5. For this purpose we take a somewhat involved test problem that will allow us to demonstrate our numerical experience with the direct solver – the goal is to take a complex-valued, non-normal matrix $L\in\mathbb{C}^{N\times N}$, for which we can directly control the non-normality as well as its definitness.
We do that by combining an HPD part and a skew-Hermitian part, scaling each to our liking. We assemble $W,X\in\mathbb{C}^{N\times N}$ by drawing the real and imaginary parts of the matrices from the distribution $\mathcal{N}(0,1)$, independently for each entry, and set $H=W^{\ast}W+\eta I$ and $K=\tfrac{1}{2}(X-X^{\ast})$. Hence, with high probability, $H$ is HPD for $\eta>0$ and hermitian, indefinite for $\eta<0$; on the other hand $K$ is skew-Hermitian. Taking their linear combination as
@@
with some $\gamma>0$, we obtain (with high-probability) a non-normal matrix $L$. Analogously, we also generate a random, complex-valued right-hand side vector and solve the resulting system as outlined in Section 5, with two GMRES iterations used at each level, continuing the recursion till the coarse space is a scalar problem.
In Figure 1, we illustrate some properties of $L$ and $M^{-1}L$ when $L$ is definite ($\eta=\gamma=10$) or indefinite ($\eta=-10,\gamma=10$); in both cases $L$ is well-conditioned (condition number $\leq 10^{2}$). First, we notice the numerically perfect spectral clustering [^6] The clustering is on the level $n^{2m+1}\varepsilon_{mach}$, which aligns with the rounding error level due to the assembly of $M^{-1}L=I-E^{(s,\mathrm{full})}E^{(c)}E^{(s,\mathrm{full})}$. when using Algorithm 2 combined with Corollary 4.5, precisely as predicted, regardless of the definitness of $L$. Next, if $L$ is definite, then the preconditioned system is also well-conditioned even for larger number of pre- and post-smoothings. However, when $L$ becomes indefinite, the preconditioned system becomes increasingly ill-conditioned as the number of pre- and post-smoothings grows. For $m=6$ we obtain $\kappa\approx 10^{8}$ for $L$ indefinite, compared to $\kappa\approx 10^{4}$ for $L$ definite. This sounds dangerous and it is – left preconditioned GMRES measures the convergence by the relative preconditioned residual, i.e., the direct solver may deliver the correct solution to a wrong problem. For $m=3$, i.e., the bottom-right-most subplot, the relative error of the resulting approximate solution becomes roughly of order $10^{-10}$ (while for $L$ definite, it stays at the level of $10^{-14}$).
Consider $L$ first with $\eta$ and then $\eta$, we observe this gap in conditioning of the preconditioned system to shrink (or widen) as $\eta\rightarrow 0$ (or $\eta\rightarrow+\infty$) but grows steadily as we increase $m$. We note that from practical viewpoint, the latter is not too pressing as we are usually not interested in running a larger number of pre- and post-smoothing. On the other hand, indefinite problems are of practical importance and there we advice caution as we believe that further understanding is necessary, especially for practical adaptations of Algorithm 2.
@@ 6.2 Sparsity of prolongation and restriction operators
With most multigrid methods we expect that the application of the matrices $S^{-1},P,R$ and $M_{0}^{-1}$ is cheap. For $P,R$ this classically follows from their structural sparsity while for $S^{-1}$ this is often achieved by construction, e.g., taking point-Jacobi or block-Jacobi (with small blocksize) smoothers; $M_{0}^{-1}$ then should inherit these properties by recursion. Using Algorithm 2 with the choices of operators as in Corollary 4.5 directly, these practical requirements will often be violated. In this section we consider adaptations that still outperform the widely used choices for a simple model problem.
@@
Applying Corollary 4.5 directly is possible but the prolongation and restriction operators then require the construction of $A^{-1}B$ (and its hermitian conjugate), which is a dense matrix, posing a significant bottleneck to the efficiency of the multilevel method.
One common practice addressing this issue is to reorder the DoFs, here using the red-black reordering. Taking $n=N/2$, both $A$ and $D$ become diagonal while $B$ and $C$ inherit the sparsity of $L$ (in our case having four nonzero entries per row/column). This way the prolongation and restriction operators are much more efficient and we can use Algorithm 2 with Corollary 4.5 and GMRES acceleration from Section 5 to obtain a direct solver.
Another choice, used more commonly for more challenging problems, is to approximate expensive operators by means of tensor lifting [hemker2004fourier], i.e., the idea that if the differential operator has a “tensor structure”, e.g., it is a tensor product or sum, then we can use this algebraically and approximate the ideal operators by the tensor product or sum of their lower-dimensional discretizations. We demonstrate this heuristic on the model problem above. First, we red-black reorder the DoFs of $L_{1D}$ so that
@@
and $A_{1D},D_{1D}\in\mathbb{R}^{\sqrt{N}/2\times\sqrt{N}/2}$ become diagonal. Then, we apply the algebraic construction of Corollary 4.5 to $L_{1D}$ based on its above blocking and obtain $P_{1D},R_{1D},S_{1D}^{-1}$. The approximate prolongation, restriction and smoothing operators for $L$ then follow as
@@
Clearly, the results of previous sections do not apply but due to the 1D reordering we can draw analogous conclusions for the sparsity as in the case of the 2D red-black reordering for this model problem, i.e., we get a diagonal smoother and sparse prolongation and restriction operators, with the same number of nonzeros per row/column. The choice of the smoothing parameters $\alpha$ is usually done heuristically, based on considering the smoother alone as in [ErnstGander2013], leading to the common choice $\alpha_{i}=2/3$ for all $i$, using LFA as in [Hackbusch1985] or based on a lifting strategy similarly to [lucerolorca2024optimization], where tedious analysis of the method in 1D yields optimal parameters and these are simply used for higher-dimensional problems. Adopting a similar mindset, we illustrate in Figure 2 that even with the choices in eq. (28) instead of eq. (2) our parameters $\alpha_{1},\dotsc,\alpha_{m}$ perform near-optimal compared to other choices of $\alpha$. This suggests that the algebraic spectral optimization of Corollary 4.5 extends approximately to higher–dimensional tensor–product discretizations while preserving sparse restriction and prolongation operators.
Figure 2: Spectral radius $\rho(E)$ as a function of the number of smoothing steps $m$ for the two–dimensional Laplacian system on a $16\times 16$ interior finite–difference grid preconditioned with Algorithm 2. The right–hand side is taken as $\mathbf{b}\equiv\mathbf{1}$ so that the underlying PDE solution is smooth. Gray curves: constant $\alpha\in\{0.1,\dots,1.0\}$; dashed: constant $\alpha=2/3$; thick curve with filled markers: smoothing parameters given by Corollary 4.5.
We also want to note that using either of these restriction/prolongation operators, the coarse matrix $M_{0}$ becomes denser than $L$ itself, e.g., in our case $M_{0}$ has the sparsity pattern of the 9-point stencil and so neither of these techniques give a win-all answer (in terms of efficiency of the multilevel method). The point is that the results of Corollary 4.5 on the optimal parameters can be used also in the tensor-lifting setting. Further testing and justification of these observations remain an active areas of research for us.
@@ 6.3 Discontinuous Galerkin 1D
The clustering result of Corollary 4.5 is purely algebraic but similar work has been recently explored also from the geometrical point of view in [LuceroLorcaGander2022] for the discontinuous Galerkin method applied to a simple test problem, where the author find optimal parameters within a family of prolongation and restriction operators. In [LuceroLorcaGander2022], the prolongation operator was designed under strict locality and symmetry constraints: it acted within each DG element and its sparsity pattern ensured that each fine-grid degree of freedom depends on a single coarse element. The interpolation was therefore purely local, and the adjustable parameters controlling its behavior were the discontinuity parameter $c$ combined with the penalty parameter $\delta_{0}$ and the relaxation parameter $\alpha$. Under these constraints, the authors showed there exists a unique combination of $(\alpha,c,\delta_{0})$ that produces a perfectly clustered spectrum.
@@
considering two admissible 11D red-black reorderingselement-wise (grouping DoFs by alternating elements) and interface-wise (alternating traces across interfaces). These lead to different local algebraic stencils for $P$ and for the Jacobi smoother $S^{-1}$. In all cases, the smoothing relaxations are the $\{\alpha_{i}\}_{i=1}^{m}$ from Corollary 4.5.
@@
The definition of $R_{\mathrm{RB}}$ and $M_{0}$ follows directly from the symmetric two-level formulation introduced in Section 2; it is not an additional assumption but a necessary consequence of the formulation. In particular, the theory developed in Section 3 applies for an arbitrary number of smoothing steps $m$, with the relaxation parameters $\{\alpha_{j}\}_{j=1}^{m}$ chosen as in Corollary 4.5. Obtaining such a closed-form description of the spectrum for general $m$ would be extremely cumbersome with a classical Local Fourier Analysis approach, which requires a mode-by-mode treatment and becomes rapidly intractable as the number of smoothing steps increases.

### v3

@@ 6.1 Conditioning of the preconditioned system
We start by numerically testing the direct solver version of Algorithm 2 with Theorem 4.5. For this purpose we take a somewhat involved test problem that will allow us to demonstrate our numerical experience with the direct solver – the goal is to find a complex-valued, non-normal matrix $L\in\mathbb{C}^{N\times N}$, for which we can directly control the non-normality as well as its definiteness.
We do that by combining an HPD part and a skew-Hermitian part, scaling each to our liking. We assemble $W,X\in\mathbb{C}^{N\times N}$ by drawing the real and imaginary parts of the matrices from the distribution $\mathcal{N}(0,1)$, independently for each entry, and set $H=W^{\ast}W+\eta I$ and $K=\tfrac{1}{2}(X-X^{\ast})$. Hence, with high probability, $H$ is HPD for $\eta>0$ and Hermitian indefinite for $\eta<0$; on the other hand $K$ is skew-Hermitian. Taking their linear combination as
@@
with some $\gamma>0$, we obtain (with high-probability) a non-normal matrix $L$. Analogously, we also generate a random, complex-valued right-hand side vector and solve the resulting system as outlined in Section 5, with two GMRES iterations used at each level, continuing the recursion until the coarse space is a scalar problem.
In Figure 1, we illustrate some properties of $L$ and $M^{-1}L$ when $L$ is definite ($\eta=\gamma=10$) or indefinite ($\eta=-10,\gamma=10$); in both cases $L$ is well-conditioned (condition number $\leq 10^{2}$). First, we notice the numerically perfect spectral clustering [^5] The clustering is on the level $n^{2m+1}\varepsilon_{mach}$, which aligns with the rounding error level due to the assembly of $M^{-1}L=I-E^{(s,\mathrm{full})}E^{(c)}E^{(s,\mathrm{full})}$. when using Algorithm 2 combined with Theorem 4.5, precisely as predicted, regardless of the definitness of $L$. Next, if $L$ is definite, then the preconditioned system is also well-conditioned even for larger numbers of pre- and post-smoothings. However, when $L$ becomes indefinite, the preconditioned system becomes increasingly ill-conditioned as the number of pre- and post-smoothings grows. For $m=6$ we obtain $\kappa\approx 10^{8}$ for $L$ indefinite, compared to $\kappa\approx 10^{4}$ for $L$ definite. This is dangerous – left preconditioned GMRES measures the convergence by the relative preconditioned residual, i.e., the direct solver may deliver the correct solution to a wrong problem. For $m=3$, i.e., the bottom-rightmost subplot, the relative error of the resulting approximate solution becomes roughly of order $10^{-10}$ (while for $L$ definite, it stays at the level of $10^{-14}$).
We observe that the gap in conditioning of the preconditioned system shrinks as $\eta\rightarrow 0^{\pm}$ and widens as $\eta\rightarrow\pm\infty$. The conditioning grows steadily as we increase $m$. We note that from practical viewpoint, the latter is not too pressing as we are usually not interested in running a larger number of pre- and post-smoothings. On the other hand, indefinite problems are of practical importance and there we advise caution as we believe that further understanding is necessary, especially for practical adaptations of Algorithm 2.
@@ 6.2 Sparsity of prolongation and restriction operators
With most multigrid methods we expect that the application of the matrices $S^{-1},P,R$ and $M_{0}^{-1}$ is cheap. For $P,R$ this classically follows from their structural sparsity while for $S^{-1}$ this is often achieved by construction, e.g., taking point-Jacobi or block-Jacobi (with small blocksize) smoothers; $M_{0}^{-1}$ should then inherit these properties by recursion. Using Algorithm 2 with the choices of operators as in Theorem 4.5 directly, these practical requirements will often be violated. In this section we consider adaptations that still outperform the widely used choices for a simple model problem.
@@
Applying Theorem 4.5 directly is possible but the prolongation and restriction operators then require the construction of $A^{-1}B$ (and its Hermitian conjugate), which is a dense matrix, posing a significant bottleneck to the efficiency of the multilevel method.
One common practice addressing this issue is to reorder the DoFs, here using the red-black reordering. Taking $n=N/2$, both $A$ and $D$ become diagonal while $B$ and $C$ inherit the sparsity of $L$ (in our case having four nonzero entries per row/column). This way the prolongation and restriction operators are much more efficient and we can use Algorithm 2 with Theorem 4.5 and GMRES acceleration from Section 5 to obtain a direct solver.
Another choice, used more commonly for more challenging problems, is to approximate expensive operators by means of tensor lifting [hemker2004fourier], i.e., if the differential operator is a tensor product or sum, then we can use this algebraically and approximate the ideal operators by the tensor product or sum of their lower-dimensional discretizations. We demonstrate this heuristic on the model problem above. First, we red-black reorder the DoFs of $L_{1D}$ so that
@@
and $A_{1D},D_{1D}\in\mathbb{R}^{\sqrt{N}/2\times\sqrt{N}/2}$ become diagonal. Then, we apply the algebraic construction of Theorem 4.5 to $L_{1D}$ based on its above blocking and obtain $P_{1D},R_{1D},S_{1D}^{-1}$. The approximate prolongation, restriction and smoothing operators for $L$ then follow as
@@
Clearly, the results of previous sections do not apply but due to the 1D reordering we can draw analogous conclusions for the sparsity as in the case of the 2D red-black reordering for this model problem, i.e., we get a diagonal smoother and sparse prolongation and restriction operators, with the same number of nonzeros per row/column. The choice of the smoothing parameters $\alpha$ is usually done heuristically, based on considering the smoother alone as in [ErnstGander2013], leading to the common choice $\alpha_{i}=2/3$ for all $i$, using LFA as in [Hackbusch1985] or based on a lifting strategy similarly to [lucerolorca2024optimization], where tedious analysis of the method in 1D yields optimal parameters and these are simply used for higher-dimensional problems. Adopting a similar mindset, we illustrate in Figure 2 that even with the choices in eq. (28) instead of eq. (2) our parameters $\alpha_{1},\dotsc,\alpha_{m}$ perform near-optimal compared to other choices of $\alpha$. This suggests that the algebraic spectral optimization of Theorem 4.5 extends approximately to higher–dimensional tensor–product discretizations while preserving sparse restriction and prolongation operators.
Figure 2: Spectral radius $\rho(E)$ as a function of the number of smoothing steps $m$ for the two–dimensional Laplacian system on a $16\times 16$ interior finite–difference grid preconditioned with Algorithm 2. The right–hand side is taken as $\mathbf{b}\equiv\mathbf{1}$ so that the underlying PDE solution is smooth. Gray curves: constant $\alpha\in\{0.1,\dots,1.0\}$; dashed: constant $\alpha=2/3$; thick curve with filled markers: smoothing parameters given by Theorem 4.5.
We also want to note that using either of these restriction/prolongation operators, the coarse matrix $M_{0}$ becomes denser than $L$ itself. In our case, $M_{0}$ has the sparsity pattern of the 9-point stencil and so neither of these techniques come without drawbacks in terms of efficiency of the multilevel method.
@@ 6.3 Discontinuous Galerkin 1D
The clustering result of Theorem 4.5 is purely algebraic but similar work has been recently explored also from the geometrical point of view in [LuceroLorcaGander2022] for the discontinuous Galerkin method applied to a simple test problem, where the authors find optimal parameters within a family of prolongation and restriction operators. In [LuceroLorcaGander2022], the prolongation operator was designed under strict locality and symmetry constraints: it acted within each DG element and its sparsity pattern ensured that each fine-grid degree of freedom depends on a single coarse element. The interpolation was therefore purely local, and the adjustable parameters controlling its behavior were the discontinuity parameter $c$ combined with the penalty parameter $\delta_{0}$ and the relaxation parameter $\alpha$. Under these constraints, the authors showed there exists a unique combination of $(\alpha,c,\delta_{0})$ that produces a perfectly clustered spectrum.
@@
considering two admissible 11D red-black reorderingselement-wise (grouping DoFs by alternating elements) and interface-wise (alternating traces across interfaces). These lead to different local algebraic stencils for $P$ and for the Jacobi smoother $S^{-1}$. In all cases, the smoothing relaxations are the $\{\alpha_{i}\}_{i=1}^{m}$ from Theorem 4.5.
@@
The definition of $R_{\mathrm{RB}}$ and $M_{0}$ follows directly from the symmetric two-level formulation introduced in Section 2; it is not an additional assumption but a necessary consequence of the formulation. In particular, the theory developed in Section 3 applies for an arbitrary number of smoothing steps $m$, with the relaxation parameters $\{\alpha_{j}\}_{j=1}^{m}$ chosen as in Theorem 4.5. Obtaining such a closed-form description of the spectrum for general $m$ would be extremely cumbersome with a classical Local Fourier Analysis approach, which requires a mode-by-mode treatment and becomes rapidly intractable as the number of smoothing steps increases.